OpenStudy (anonymous):

(x+6)(x-7)(x+8(x+9)=0 Solve the equation

OpenStudy (campbell_st):

well if the question is (x +6)(x-7)(x +8)(x+9) = 0 if any binomial factor(s) equal zero then the whole equation will equal zero. so you need to solve each binomial factor x + 6 = 0 x -7 = 0 x + 8 = 0 and x + 9 = 0 so there will be 4 unique solutions.